OpenStudy (anonymous):

{4x - 2y = -6 {12x - 3y = -15 Solve by either substitution or elimination.

OpenStudy (anonymous):

@sangya21 help me

OpenStudy (anonymous):

By elimination: -Multiply first eq. by 3 and then subtract it from second eq. By substitution: - 2(2x -y) = -6 2x-y = -3 2x+3 = y now put the value of y in second eq to find x then use that value of x to find y